### v1

@@ 3.1 Multi-task learning via robust regularized clustering
Existing MTL methods based on clustering have not taken into account the presence of outlier tasks. For instance, CVXMTL estimates the regression coefficients uniformly close to the corresponding estimated centroids, which ignores the nature of outlier tasks that have a large task-specific characteristic or no relationship to other tasks. Some robust MTL methods (Chen et al., 2011; Gong et al., 2012) have attempted to address the issues of outlier tasks by introducing outlier parameters and selecting them using group lasso regularization. However, group lasso (Yuan and Lin, 2006) limits the value of the outlier parameters, which may not adequately represent their nature.
@@ 6 Application to real datasets
Table 3: AUC and NMSE for landmine data and school data in 100 repetions
@@
MTLRRC (GS) 0.759 (0.022) $\bm{0.815}$ (0.029)
MTLRRC (GM) 0.758 (0.020) 0.816 (0.030)
MTLCVX $\bm{0.761}$ (0.024) 0.822 (0.069)
MTLK 0.755 (0.021) 0.816 (0.039)
(a) Mean of estimated $\widehat{W}$ for 100 repetitions in landmine data
(b) Mean of estimated $\widehat{U}$ for 100 repetitions in landmine data
(c) Mean of estimated $\widehat{O}$ for 100 repetitions in landmine data
Figure 2: The mean of the estimated value of parameters in MTLRRC with group SCAD in 100 repetitions for the landmine data.
(a) Mean of estimated $\widehat{W}$ for 100 repetitions in school data
(b) Mean of estimated $\widehat{U}$ for 100 repetitions in school data
(c) Mean of estimated $\widehat{O}$ for 100 repetitions in school data
Figure 3: The mean of the estimated value of parameters in MTLRRC with group SCAD in 100 repetitions for the school data.
@@
We compared MTLRRC with CVXMTL and MTLK. First, we split the samples in each task into $60\%$ for the train, $20\%$ for the validation, and $20\%$ for the test. The regularization parameters are determined by the validation data. For the non-convex penalties in MTLRRC, we fixed $\gamma=3$ for the group MCP and $\gamma=3.7$ for the group SCAD. For application to the school data, we fixed $\lambda_{2}=1000$ in MTLRRC and MTLCVX so that there is only one estimated cluster, because the school data has been considered rather homogeneous in some studies (Bakker and Heskes (2003); Evgeniou et al. (2005)). For the evaluation, we calculated NMSE and AUC for school data and landmine data, respectively. These values were calculated 100 times with the random splitting of the dataset.
Table 3 shows the mean and standard deviation in 100 repetitions. From this table, we observe that MTLRRC gives a relatively smaller NMSE than other methods for the school data, while MTLCVX is superior to other methods for the landmine data.
Next, we calculated the mean of estimated parameters $\widehat{W},\widehat{U}$, and $\widehat{O}$ in MTLRCC with group SCAD. Figures 2 and 3 show the mean of estimated parameter values for the landmine data and the school data, respectively. The vertical axis is the index of tasks and the horizontal axis is the index of features. Each color shows the mean of the estimated parameters. Figures 4 and 5 are bar plots that show the ratios of each task detected as an outlier task. Here, note that tasks that have never been detected are removed from the bar plots.
For the landmine data, Figure 2(2) suggests the presence of two clusters, which is consistent with the fact that tasks 1–15 and 16–29 are obtained from regions corresponding to different surface conditions. For outliers detection in Figure 2(2), because the $\bm{o}_{m}$ for $10$-th task has a relatively larger or smaller value than other tasks and the task is only a task detected as an outlier with a ratio greater than 0.3 from Figure 4, it may indicate that the task is a potential outlier task. The estimated regression coefficients for the $10$-th task in Figure 2(2) show a similarity to other tasks within the same estimated cluster. These results may suggest the underlying structure among tasks in the landmine data is rather Case 1 than Case 2 in Section 5. Hence, it is not surprising that MTLRRC is inferior to MTLCVX in terms of prediction accuracy. Furthermore, from Figure 4, we observe that only one task in tasks 16–29 has an outlier task ratio of more than 0.1, while eight tasks in tasks 1–15 have. This result suggests that the cluster composed of tasks 1–15 may have relatively large variability. This may provide some insight into the structure within the cluster.
For the school data, by setting $\lambda_{2}=1000$, we obtained the homogeneous pattern in Figure 3(3). From Figure 5, 2nd, 28th, 66th, 76th, and 83rd tasks were detected with a rate greater than 0.4. In addition, these tasks show larger values in $\bm{o}_{m}$ than other tasks from Figure 3(3). For the regression coefficients in Figure 3(3), although these tasks share many characteristics with other tasks, some regression coefficients have different characteristics. For example, the regression coefficients for $76$th and $83$rd tasks of $16$th feature have, respectively, relatively large and small values that are much different from other tasks.

### v2

@@ 3.1 Multi-task learning via robust regularized clustering
Existing MTL methods based on clustering have not taken into account the presence of outlier tasks. For instance, MTLCVX estimates the regression coefficients uniformly close to the corresponding estimated centroids, which ignores the nature of outlier tasks that have a large task-specific characteristic or no relationship to other tasks. Some robust MTL methods (Chen et al., 2011; Gong et al., 2012) have attempted to address the issues of outlier tasks by introducing outlier parameters and selecting them using group lasso regularization. However, group lasso (Yuan and Lin, 2006) limits the value of the outlier parameters, which may not adequately represent their nature.
@@ 6 Application to real datasets
Table 3: AUC and NMSE for landmine data and school data in 100 repetitions
@@
MTLRRC (GS) $\bm{0.764}$ (0.021) 0.853 (0.058)
[…]
MTLRRC (GM) 0.760 (0.026) 0.852 (0.053)
[…]
MTLCVX 0.760 (0.022) 0.847 (0.048)
MTLK 0.756 (0.023) 0.847 (0.048)
(a) Mean of estimated $\widehat{W}$ for 100 repetitions in the landmine data
(b) Mean of estimated $\widehat{U}$ for 100 repetitions in the landmine data
(c) Mean of estimated $\widehat{O}$ for 100 repetitions in the landmine data
[…]
(a) Mean of estimated $\widehat{W}$ for 100 repetitions in the school data
(b) Mean of estimated $\widehat{U}$ for 100 repetitions in the school data
(c) Mean of estimated $\widehat{O}$ for 100 repetitions in the school data
[…]
@@
[…]
For the landmine data, Figure 2(2) suggests the presence of two clusters, which is consistent with the fact that tasks 1–15 and 16–29 are obtained from regions corresponding to different surface conditions. For outliers detection in Figure 2(2), because the $10$-th task has a relatively larger or smaller value than other tasks and the task is a task detected as an outlier with a ratio greater than 0.3 from Figure 4, it may indicate that the task is a potential outlier task. Furthermore, the $9$-th task also has a relatively high detection ratio that is greater than 0.3, although the mean of $\widehat{\bm{o}}_{m}$ is not as clear as the $10$-th task from Figure 2(2). The estimated regression coefficients for the $9$-th and $10$-th tasks in Figure 2(2) show a similarity to other tasks within the same estimated cluster. These results may suggest the underlying structure among tasks in the landmine data is rather Case 1 than Case 2 in Section 5. Furthermore, from Figure 4, we observe that only two tasks in tasks 16–29 have an outlier task ratio of more than 0.1, while 13 tasks in tasks 1–15 have. This result suggests that the cluster composed of tasks 1–15 may have relatively large variability. This may provide some insight into the structure concerning sub-groups within the cluster.
For the school data, we obtained the homogeneous pattern in Figure 3(3). The school data have been considered rather homogeneous in some studies (Bakker and Heskes (2003); Evgeniou et al. (2005)). Thus, this result would be reasonable and shows that MTLRRC can consistently estimate cluster structure even when the true number of underlying clusters is one. From Figure 5, the 2nd, 28th, 66th, 76th, and 83rd tasks were detected with a rate greater than 0.3. In addition, these tasks show larger values in $\bm{o}_{m}$ than other tasks from Figure 3(3). For the regression coefficients in Figure 3(3), although these tasks share many characteristics with other tasks, some regression coefficients have different characteristics. For example, the regression coefficients for $76$th and $83$rd tasks of $16$th feature have, respectively, relatively large and small values that are much different from other tasks.